A researcher examined two factors that might predict self-esteem in college students: their grades in a Stats 2 class, and the number of followers they have on their Instagram account. She surveyed some 53 students and obtained the following values: SDY = 20, MSM of 2000, and a significant F value. She would really love it if her model generalized outside of her original sample. What is her Adjusted R2, and does the model still account for a significant amount of variance in other samples (Make sure to show your obtained F value, along with the degrees of freedom, to support your answer)?
Answer #1

Here we have N = 53, Standard deviation = 20, MSM = 2000 and k =2.

SSM = Sum of squares For Model

SSE = Sum of squares For Error

SST = Total Sum of squares

MSM = mean sum of squares for model

MSE = mean sum of squares for error

SDY2= total variation in the response y

Sources SS df MS F
Model SSM k-1 SSM/k-1 MSM/MSE
Error SSE N-k SSE/N-k
Total SST N-1

Now k = 2 implies k-1 = 2-1 = 1.

Also, N-k = 53 - 2 = 51

Now, MSM = 2000 implies SSMSSM k-1 MSM =

200 SS

SSM2000

Now, SSM+ SSE = SST implies 2000 + SSE = SST

Also, Total variance = SD2 = 400

Hence 2000 + SSE = 400 which implies SSE = 1600.

MSE =1600/51 = 31.37

F-value = MSM / MSE = 2000/47.06 = 63.75

ANOVA TABLE:

Sources SS df MS F
Model 2000 1 2000 63.75
Error 1600 51 31.37
Total 400 52 7.69

Adjusted R-squared value = MSE 31.37 MST ,69-=-3,08